1. If hydrogen-3 has a half-life of 12.3 years, how many years will it take for 109.9 mg 3H to decay to 0.013 mg 3H?
Answer #1

A/c to N = N0 (1/2)n

N/N0 = (1/2)n

No of halflifes (n) to reduce 109.9 to 0.013 = 13

Total time = n x half life

   = 12.3 x 13 = 159.9 yrs